Three times the greater of three consecutive even integers exceeds twice the least by 38. What are the integers
Answer by user_dude2008(1862)

3(x+4)=2x+38
3x+12=2x+38
3x-2x=38-12
x=26 <--------first number
x+2=28 <--------second number
x+4=30 <--------third number